# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 5.1 Our novelties (a short list)
Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near optimal level.
@@ 5.2 Related works
The pioneering paper [76] proposed a novel algorithm for approximation of an absolutely largest zero of a real-rooted black box polynomial, estimated the computational precision of that algorithm, and extended it to approximation of an absolutely largest eigenvalue of a symmetric matrix. From that paper we adopt the model of computation, extension to approximation of matrix eigenvalues, and some patterns for estimation of the computational precision of root-finders, but otherwise use distinct approach and techniques.
Namely, instead of Newton’s iterations in [76], we accelerate the classical subdivision root-finders, based on e/i tests. Its closest variations in [128, 101, 23, 24] still use inferior e/i tests, for which we proposed two fast novel approaches – based on approximation of Cauchy integrals and on root-lifting; these tests run much faster than the tests in the cited papers and, unlike the latter tests, can be applied to a black box polynomial. The papers [128, 101, 23, 24] apply functional iterations (such as Newton’s and Schröders) to accelerate convergence of subdivision process but we apply our novel techniques of disc compression instead.
@@
Estimating the bit operation complexity of our root-finders we use [134, Eqn. (9.4)] and [81, Thm. 2] for multipoint polynomial evaluation. [^8] The ingenious algorithm of Moroz supporting his [81, Thm. 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$. To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$. At this point, application of [68, Alg. 5.3 and Thm. 3.9] yields new record fast algorithm for multipoint approximation of a polynomial given with its coefficients. The root-finders of Moroz in [81] and Imbach and Moroz in [59] use $\tilde{O}(bd)$ bit operations for polynomials given with their coefficients and having only well-conditioned zeros. In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [14, 30, 22], the upper bound on the condition numbers of all zeros of $p(x)$ in [81, 59] is a little stronger restriction than the bound on the cluster sizes of Def. 24.1, under which [96, 103] proved the same complexity bound. Unlike our present root-finders, the root-finders of [81, 59] are not efficient for the worst case inputs and extensively use the coefficients.
5.3 Implementation: State of the Art
@@
The near-optimal but quite involved algorithm of [96, 103] has never been implemented, while the slower algorithm of [134] has been implemented by Xavier Gourdon in [46] for the Magma and PARI/GP computer algebra systems. By now it is superseded by implementation of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (in Sec. 26.4 we list some competitive implementations of Newton’s iterations).
@@
We present algorithms that are rather simple for implementation and promise to compete for user’s choice. Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder [64] has enabled its noticeable acceleration (see [60]). The more advanced publicly available implementation in [63] “relies on novel ideas and techniques" from [112] (see [63, Sec. 1.2]). According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [64], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^9] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [96, 103] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots. Thus already implementation in [64] of subdivision iterations, slightly outperformed MPSolve for root-finding in a domain $\mathbb{D}$ containing a small number of roots although is by far inferior for the approximation of all $d$ roots. and tended to supersede MPSolve more and more significantly as the degree $d$ of $p$ grew large.
@@
There is still a large room for further improvement because the root-finder of [63] inherits some earlier sub-algorithms from [23, 64], which are slower by order of magnitude than our current ones.
@@
Next we comment on some promising directions to further improvement of [63], but also notice that adjustment of our near-optimal subdivision root-finder in Remark 2.1 to approximation of real and near-real roots and consequently of matrix eigenvalues should be at least competitive with the current user’s choice [70] for real root-finding as well as with its significant improvement in [62]; moreover, it more efficiently counters the impact of the input and rounding errors.
5.4 Further progress and promise
(A) Acceleration of e/i tests. E/i tests are basic ingredient of subdivision root-finders. [60, 63] used the Cauchy-based e/i tests of our Part II and their heuristic acceleration. It is still a challenge to compare their empirical performance with that of the formally supported e/i tests of our Part III and Sec. 23.4, based on root-lifting. Furthermore, we can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [146]), efficiently implemented in Boost C++ libraries.
(B) Acceleration of the compression of a disc. We assume that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the external $d-m$ roots, and we compress that disc towards the minimal disc containing the same $m$ roots. We reduce this task essentially to (i) computing a point in or near the convex hull of the root set in $D(0,1)$ and (ii) approximation of the $m$th smallest root radius from a given point in $D(0,1)$. For task (i) we dramatically accelerate our algorithm of [60, 63] because our algorithms of Secs. 23.2 and 23.3 for task (i), using Schröder’s iterations, deterministic, and formally supported, run dramatically faster than those of [60, 63], based on approximation of the average of all roots in $D(0,1)$. For task (ii) we can apply our Las Vegas randomized algorithms of Secs. 8 and 9 combined or the alternative of [60, 63], based on the algorithm of our Sec. 15.1, which reduces this task to performing a constant number of e/i tests. Both approaches support the same asymptotic cost bound up to polylog factors, but fro $m\gg\log^{2}(d)$ the second approach is faster by a factor of $m/\log^{2}(d)$ where it uses fast heuristic e/i tests. Such an improvement, however, may be lost in practice because of the impact of polylog factors. Implementation of compression in [60, 63] is slowed down with verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our Las Vegas root-finders.
@@ 5.4 Further progress and promise
(C) Fast approximation of the eigenvalues of a matrix. [20] explored reduction of approximation of the eigenvalues of a matrix to root-finding by means of Ehrlich’s iterations, reduced to recursive computations of NIR$(x)$, and showed benefits of this approach. We can readily extend this reduction and its benefits by using subdivision iterations instead of Ehrlich’s. Furthermore, win the case of subdivision we obtain additional benefits, e.g., our algorithms are much faster for approximation of a small number of eigenvalues that lie in a fixed domain and unlike Ehrlich’s iterations isolate eigenvalues, rather than just approximate them.
@@
(D) Acceleration of real and near real root-finding. In Remark 2.1 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors and can be readily extended to approximation of the eigenvalues of a matrix.
@@ 5.5 Organization of the rest of the paper
In Parts II and III we devise and analyze our two alternative e/i tests. Deterministic e/i test of Part III is as fast as the randomized one of Part II. We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost, and in Part III we yield acceleration by a factor of $d$ under the Random Root Model. The bit-operation complexity bounds of Parts II and III are inferior by a factor of $b$ to that of [96, 97, 103], but we match it in Part IV by means of our acceleration of the algorithms of both Parts II and III based on compression of a disc without losing its roots. In Part V we cover some important root-finding subjects that involve no substitution iterations, namely, factorization, root isolation, and functional iterations. In Secs. 24 and 25 we outline the fundamental results of Schönhage from [134] on polynomial factorization, its links to root-finding and root-isolation, and some relevant bit-complexity estimates, including lower bounds. In Sec. 26 we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and their acceleration by means of incorporation of Fast Multipole Algorithm (FMM) and fast multipoint polynomial evaluation as well as their initialization both where the coefficients of $p$ are unknown or given to us.
5.6 Independence of Part III from Part II
In Parts II an III we solve Problems 0, 1, 0, and 1 in two different ways. In Part II we rely on approximation of a Cauchy integral associated with root-counting in a disc on the complex plane. In Part III we rely on lifting the roots to a fixed power $h=O(\log(d))$. In addition, in Part III, unlike Part II, we devise a faster root-finder under a random root model. We make our presentation in Part III independent of Part II; in particular we include again some definitions, figures, and theorems.

### v37

@@ 5.1 Our novelties (a short list)
Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near-optimal level.
@@ 5.2 Related works
From the pioneering paper [59] we adopt the model of computation, extension of a black box polynomial root-finder to approximation of matrix eigenvalues, and the approach to estimation of the computational precision of root-finders, but otherwise we use distinct approach and techniques.
@@
Estimating the bit operation complexity of our root-finders we use [106, Eqn. (9.4)] and [64, Thm. 2] for multipoint polynomial evaluation. [^11] The ingenious algorithm of Moroz supporting his [64, Thm. 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$. To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$. At this point, application of [52, Alg. 5.3 and Thm. 3.9] yields new record fast algorithm for multipoint approximation of a polynomial given with its coefficients. Ignoring Kirinnis’s paper [52], Moroz cited a rediscovery of his result by Kobel and Sagraloff, made two decades later.
Instead of applying high order Newton’s iterations of [59], we accelerate classical subdivision root-finders. The closest earlier variations in [100, 82, 16, 17] use mach slower e/i tests, involving coefficients of $p$. Two our fast novel e/i tests – based on root-lifting and approximation of Cauchy integrals, respectively, work for a black box polynomial.
The algorithms of [100, 82] dramatically accelerate convergence of subdivision process by means of applying Newton’s or Schröder’s iterations (see Fig. 3). They converge with superlinear rate to the convex hull of all roots lying in $C$. The diameter of the minimal covering disc, however, can be too large, and then the papers [100, 82] apply subdivision iterations again. The resulting root-finders recursively intertwine subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder.
[100, 82] have only estimated the number of ops but not of the bit operations involved. [16, 17] estimated the bit operation complexity in terms of the unknown minimal distance between the unknown roots. Specified in terms of $b$ and $d$, the estimates turned out to be inferior to ours by a factor of $d$ (see Remark 4.2 and Sec. 10.4).
Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [106]. It first splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively. Then it recursively splits the polynomial $p(x)$ into factors $f=f(x)$ and $g=g(x)$. The resulting algorithm of [106] is slower than ours by a factor of $d$. It has been accelerated to a near-optimal level in [78, 84], although that root-finder is much involved and extensively uses the coefficients of $p$, like the root-finders of [106, 100, 82, 16, 17].
[64, 44] presented efficient root-finders for polynomials having only well-conditioned zeros, but not for worst case polynomials $p(x)$. Furthermore, the algorithms of [64, 44] extensively use the coefficients of $p(x)$.
5.3 Implementation: State of the Art (briefly)
@@
The near-optimal but quite involved algorithm of [78, 84] has never been implemented, while the slower algorithm of [106] has been implemented by Xavier Gourdon in [34] for the Magma and PARI/GP computer algebra systems. By now it is superseded by implementations of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (see some competitive implementations of Newton’s iterations in [115, 102, 68]).
@@
We present algorithms that promise to compete for user’s choice. Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder [49] has enabled its noticeable acceleration (see [45]). The more advanced publicly available implementation in [48] “relies on novel ideas and techniques" from [93] (see [48, Sec. 1.2]). According to the extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [49], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^12] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [78, 84] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots. Thus already implementation in [49] of subdivision iterations, slightly outperformed MPSolve for root-finding in a domain $\mathbb{D}$ containing a small number of roots although is by far inferior for the approximation of all $d$ roots. and superseded MPSolve more and more significantly as the degree $d$ of $p$ grew large.
@@
There is still a large room for further improvement because the root-finder of [48] inherits some earlier sub-algorithms from [16, 49], which are slower by order of magnitude than our current ones (see the next subsection).
@@
Application of our near-optimal subdivision root-finder in Remark 2.2 to approximation of real and near-real roots promises to be at least competitive with user’s current choice [54] and its improvement in [47]. It can be applied to a black box polynomial, counters the impact of input and rounding errors, and can be extended to approximation of real and near-real matrix eigenvalues.
5.4 Further progress and promises
(A) Acceleration of e/i tests. E/i tests are basic ingredient of subdivision root-finders. [45, 48] used the Cauchy-based e/i tests of our Part III and their heuristic acceleration. We can devise and test similar heuristic acceleration of the formally supported e/i tests of our Part II, based on root-lifting. Namely, we can merely apply Alg. 7.2 with an integer $q$ decreased dynamically from $m+1$, as long as the exclusion claim remains empirically correct.
(B) Acceleration of the compression of a disc. Assuming that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the $d-m$ external roots, we compress that disc towards an $\epsilon$-disc or a rigid disc, in both cases losing no roots. Algorithmic extras 1 (for compression of a disc with a small number of roots) and 2 of Sec. 21 can help strengthen this approach.
Like us, [48] reduces compression to (a) non-costly root-counting in an isolated disc, (b) computing a point $c$ lying in or near the MCD, that is, the minimal covering disc  [^13] One may use the convex hull instead of the minimal covering disc. of the root set in $D(0,1)$, and (c) approximation of the $m$th smallest root radius from $c$.
At stage (b) [48] computes a close approximation of the sum $s_{1}$ of the $m$ roots lying in the disc $D(0,1)$ and then outputs $c=\frac{s_{1}}{m}$. In Sec. 17 we greatly accelerate this algorithms in the case where the radius of the output disc is small.
At stage (c) [48] applies the algorithm of our Sec. 18, which performs a constant but fairly large number of e/i tests in a disc. One may test this approach versus our alternative of Sec. 13.4 using a single e/i test for an annulus and versus the combination of the Las Vegas randomized algorithm of Sec. 13 with BoL of Sec. 16.
Finally, implementation of compression in [45, 48] is slowed down at the stage of verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our root-finders at a dominated cost (see Sec. 3.2).
(C) Fast approximation of the eigenvalues of a matrix. Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can extend them to approximation of the eigenvalues of a matrix or a matrix polynomial by following [59] or based on the following alternative expressions:
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
where $M$ is a matrix, while $p(x)=\det(xI-M)$ (cf. (3.1)) and
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [13, Eqn. (5)]. [^14] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant.
These expressions reduce approximation of the eigenvalues of a matrix or a matrix polynomial to its inversion. One can use the matrix inversion algorithm and its bit operation cost estimates of [110] towards alternative derivation of our Main Theorem 2 for any input matrix or matrix polynomial, but the inversion can be greatly accelerated in the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
@@ 5.4 Further progress and promises
Bini and Noferini in [13] proposed and explored such a reduction of approximation of all $d$ eigenvalues to root-finding by means of Ehrlich’s iterations and showed various benefits of this approach, but we can readily combine this reduction with subdivision iterations instead of Ehrlich’s and then obtain additional benefits: our algorithms (a) are much faster for approximation of a small number of eigenvalues that lie in any fixed domain and (b) isolate the eigenvalues instead of just approximating them.
@@
(D) Acceleration of real and near-real root-finding. In Remark 2.2 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors.
@@ 5.5 Organization of the rest of the paper
In Parts II and III we devise and analyze our two e/i tests. We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost. In Part II the same randomization helps us devise fast e/i tests. In Part III we support that step with additional randomization.
In Part II we further accelerate by a factor of $d$ e/i tests and overall root-finding under the Random Root Model.
We make our presentation in Part III largely independent of Part II by reusing some definitions, results, and arguments.
In Part IV we devise and analyze our compression algorithm and incorporate it into the algorithms of Parts II and III – to accelerate them by a factor of $b/\log(b)$ and to ensure pairwise isolation of the output roots and their clusters.
In Part V we cover (i) low bounds on the bit-operation complexity of polynomial factorization, root-finding, and root isolation, (ii) acceleration with FMM (cf. Remark 3.1) and initialization of root-finding by means of functional iterations, and (iii) some additional algorithms for root radii approximation.